# Dynamic-Dark SLAM: RGB-Thermal Cooperative Robot Vision Strategy for Multi-Person Tracking in Both Well-Lit and Low-Light Scenes

### v1

@@ Abstract
In robot vision, thermal cameras have significant potential for recognizing humans even in complete darkness. However, their application to multi-person tracking (MPT) has lagged due to data scarcity and difficulties in individual identification. In this study, we propose a cooperative MPT system that utilizes co-located RGB and thermal cameras, using pseudo-annotations (bounding boxes + person IDs) to train RGB and T trackers. Evaluation experiments demonstrate that the T tracker achieves remarkable performance in both bright and dark scenes. Furthermore, results suggest that a tracker-switching approach using a binary brightness classifier is more suitable than a tracker-fusion approach for information integration. This study marks a crucial first step toward “Dynamic-Dark SLAM,” enabling effective recognition, understanding, and reconstruction of individuals, occluding objects, and traversable areas in dynamic environments, both bright and dark.
1 Introduction
@@
In this study, we propose a novel cooperative MPT system that utilizes an RGB camera and a thermal (T) camera placed in the same pose, enabling effective tracking in both bright and dark conditions. The key idea is to train a T-tracker using knowledge transfer from an RGB-tracker in bright conditions. Among all possible combinations of (1) two scene types (Bright/Dark), (2) two modality types (RGB/T), (3) three tracker types (RGB/T/RGBT tracker), we exclude those that are clearly ineffective or redundant, leaving four viable combinations, as discussed in Section 4, which we then evaluate experimentally. The results demonstrate that the T-tracker achieves remarkable performance in both bright and dark scenes. Furthermore, our findings suggest that a tracker-switching approach using a binary brightness classifier is more suitable than a tracker-fusion approach for integrating the two trackers.
This study builds on previous research on teacher-student knowledge transfer [7], multi-person tracking under occlusion [8], and the recognition, understanding, and reconstruction of people, occlusions, and traversable areas [9], effectively integrating thermal SLAM [10] and dynamic SLAM [3]. This marks a crucial first step toward “Dynamic-Dark SLAM,” enabling the effective recognition, understanding, and reconstruction of individuals, occluding objects, and traversable areas in both bright and dark dynamic environments.
@@ 2 Related Work
RGB-T fusion tracking has been studied in both non-deep approaches [16, 17, 18, 19] and deep approaches, including adaptive fusion [20], sparse fusion representation [21], cross-modal ranking [22], and graph models [23], as well as benchmark studies [24]. The approach of applying an RGB tracker [25] to the RGB-T tracking problem was explored in [26]. Our approach is based on the state-of-the-art [ByteTrack] tracker and differs in that it handles Dark scenes, where RGB trackers completely fail.
3 RGB-T Cooperative MPT
3.1 Formulation of MPT
@@ 3.3 Cooperative Strategy
$\hat{y}_{t}=f_{\theta^{\text{RGB}}\theta^{\text{T}}}(x_{t}^{\text{RGB}},x_{t}^{\text{T}})$ (5)
where $f_{\theta^{\text{RGB}}\theta^{\text{T}}}$ represents the fusion function integrating both modalities. This fusion model integrates the tracking results of the RGB teacher model and the T student model, selecting the result with the higher confidence score in each frame to achieve robust tracking. For each $i$ $(i\in[1,I^{t}])$, $y_{t}[i]=f^{\text{T}}(x_{t}^{\text{T}})$, $\tilde{y}_{t}[j]=f^{\text{RGB}}(x_{t}^{\text{T}})$, and for the corresponding pair $j$ $(j\in[1,I^{t}])$ satisfying $y_{t}^{\text{ID}}[i]=\tilde{y}_{t}^{\text{ID}}[j]$, the final fusion result is determined as follows:
@@ 4 Evaluation Experiment
Each tracking method is evaluated using Multiple Object Tracking Accuracy (MOTA) and ID F1 score (IDF1) to assess tracking performance, along with False Positive (FP), True Positive (TP), and False Negative (FN) to evaluate false detections [35]. ID Switches (IDSW) and Higher Order Tracking Accuracy (HOTA) measure trajectory reconstruction, with HOTA providing a global tracking assessment. Unlike MOTA, HOTA balances precision and recall by decomposing detection, assignment, and localization for a more detailed analysis (Table 1). These metrics gauge how effectively the proposed method tackles key challenges in Multiple Person Tracking (MPT): ID switching, detection failures, occlusion, and handling the appearance/disappearance of individuals. ID switching misassigns identities, detection failures disrupt continuity, and occlusion hides targets. Additionally, newly appearing or disappearing persons may be misidentified, reducing accuracy.
Table 1: Tracking Accuracy
@@
RGB 38.3 56.3 26 1191 1822
T 93.5 97.0 113 2942 71
FUSION 92.4 96.6 136 2942 71
RGB: $f_{\theta^{\text{RGB}}}(x_{\text{Dark}}^{\text{T}})$, T: $f_{\theta^{\text{T}}}(x_{\text{Dark}}^{\text{T}})$, FUSION: $f_{\theta^{\text{RGB}}\theta^{\text{T}}}(x_{\text{Dark}}^{\text{T}},x_{\text{Dark}}^{\text{T}})$.
Table 2: Trajectory Prediction Performance
@@
5 Conclusion
In this study, we tackled the cooperative MPT problem using both RGB and thermal (T) cameras placed in the same pose for the first time. As expected, the RGB tracker completely failed in the Dark scene’s RGB vision and also showed reduced performance in thermal vision. As a solution, we confirmed that learning the T tracker using pseudo-annotations from the RGB tracker is highly effective, achieving the best performance in both Bright and Dark scenes. We also investigated a fusion tracker that integrates both the RGB and T trackers, but surprisingly, its performance ranked second. This suggests that a cooperative approach dynamically switching between the two trackers is more effective than integration. In the future, we aim to apply the developed RGB-T cooperative MPT tracker to real-world robotic vision applications, including recognition, understanding, and reconstruction of individuals, occluders, and traversable areas.

### v3

@@ Abstract
In robot vision, thermal cameras hold great potential for recognizing humans even in complete darkness. However, their application to multi-person tracking (MPT) has been limited due to data scarcity and the inherent difficulty of distinguishing individuals. In this study, we propose a cooperative MPT system that utilizes co-located RGB and thermal cameras, where pseudo-annotations (bounding boxes and person IDs) are used to train both RGB and thermal trackers. Evaluation experiments demonstrate that the thermal tracker performs robustly in both bright and dark environments. Moreover, the results suggest that a tracker-switching strategy — guided by a binary brightness classifier — is more effective for information integration than a tracker-fusion approach. As an application example, we present an image change pattern recognition (ICPR) method, the “human-as-landmark,” which combines two key properties: the thermal recognizability of humans in dark environments and the rich landmark characteristics—appearance, geometry, and semantics—of static objects (occluders). Whereas conventional SLAM focuses on mapping static landmarks in well-lit environments, the present study takes a first step toward a new Human-Only SLAM paradigm, “D&D-SLAM,” which aims to map even dynamic landmarks in complete darkness. Additionally, this study demonstrates that knowledge transfer between thermal and depth modalities enables reliable person tracking using low-resolution 3D LiDAR data without RGB input, contributing an important advance toward cross-robot SLAM systems.
I Introduction
@@
In this study, we propose a novel cooperative MPT system that utilizes an RGB camera and a thermal (T) camera placed in the same pose, enabling effective tracking in both bright and dark conditions. The key idea is to train a T-tracker using knowledge transfer from an RGB-tracker in bright conditions. Among all possible combinations of (1) two scene types (Bright/Dark), (2) two modality types (RGB/T), (3) three tracker types (RGB/T/RGBT tracker), we exclude those that are clearly ineffective or redundant, leaving four viable combinations, as discussed in Section V, which we then evaluate experimentally. The results demonstrate that the T-tracker achieves remarkable performance in both bright and dark scenes. Furthermore, our findings suggest that a tracker-switching approach using a binary brightness classifier is more suitable than a tracker-fusion approach for integrating the two trackers.
TABLE I: Relationship to Prior Work.
Scene Type Well-Lit Low-Light
Landmark Type S D S D
conventional SLAM
Dynamic SLAM (e.g., [3])
Thermal SLAM (e.g., [7])
D&D-SLAM(Ours.)
S: Static Landmark. D: Dynamic Landmark.
As an application example, we propose D&D-SLAM, a novel SLAM framework that leverages dynamic landmarks in complete darkness, to recognize, understand, and reconstruct people, occluder objects, and traversable areas. This framework is motivated by recent advances in RGB and thermal-based human activity recognition [8]—particularly in identifying human-object interactions, such as sitting on a chair or leaning on a desk. Building on this insight, we introduce “human-as-landmark,” an image change pattern recognition (ICPR) method that exploits two key properties: (1) the thermal visibility of humans in the dark, and (2) the rich landmark features—appearance, geometry, and semantics—of static objects involved in human interaction. We experimentally demonstrate that even a simple implementation of human-as-landmark improves loop closure detection, a task where conventional methods frequently fail due to false negatives.
In addition to the RGB-T cooperative vision, this study also explores person tracking using Thermal (TIR) and Depth images without RGB input. By leveraging knowledge transfer across these different sensor modalities, the approach achieves reliable multi-person tracking even with low-resolution depth images derived from 3D LiDAR sensors. This extension broadens the applicability of cooperative vision systems and represents an important step toward cross-robot SLAM. Details and experimental results of this Thermal-Depth cooperative tracking are presented in Section V-C.
@@ II Related Work
RGB-T fusion tracking has been studied in both non-deep approaches [14, 15, 16, 17] and deep approaches, including adaptive fusion [18], sparse fusion representation [19], cross-modal ranking [20], and graph models [21], as well as benchmark studies [22]. The approach of applying an RGB tracker [23] to the RGB-T tracking problem was explored in [24]. Our approach is based on the state-of-the-art ByteTrack [25] tracker and differs in that it handles Dark scenes, where RGB trackers completely fail.
III Dynamic-Dark SLAM (D&D-SLAM)
The combination of low-light conditions and dynamic landmarks constitutes one of the most challenging scenarios for SLAM technologies. Most existing SLAM approaches assume the presence of some illumination (i.e., not complete darkness) and the presence of recognizable static landmarks within the scene [26]. Accordingly, much of the literature has focused on mapping and localization based on static landmarks.
Our approach takes an opposite yet complementary perspective. We propose Dynamic-Dark-SLAM (D&D-SLAM), a novel human-only SLAM framework that diverges from conventional bearing-range [27], range-only [28] or bearing-only [29] SLAM paradigms by using dynamic objects—specifically, humans—as landmarks in completely dark environments. Since RGB cameras are ineffective in total darkness, our method instead leverages thermal (T) cameras.
III-A Challenges
This paradigm shift introduces several unique challenges:
Human sensing suffers from high noise levels, large intra-class variation and non-rigid appearance (e.g., clothing, posture, accessories), and inter-person similarity, making re-identification difficult.
Landmark dynamics are significant because human landmarks frequently appear, disappear, and move. This movement compromises the 100% precision [30] required for loop closure, risking catastrophic mapping errors. It is noteworthy that the HO3 landmarks are intentionally transient—they may appear, disappear, or move—and are used primarily for loop closure rather than for persistent long-term map storage.
Sensor limitations include difficulty in recognizing appearance, geometry, and semantics under thermal imaging, absence of color cues, and variability in thermal signatures due to body temperature changes.
III-B Related Work and Inspirations
Our approach is inspired by two lines of prior research. First, loop-closure detection and image change detection (LCD-ICD) approaches [31] and its follow-ups [32] demonstrated that when the relative position between a landmark A and an object B remains unchanged, both can be considered static. Second, human-only SLAM works [33] and [34] treat the occlusion boundary between a human A and an occluding object B as a pseudo-static landmark feature, albeit prone to frequent false negatives.
Figure 3: Block diagram of the HO3-SLAM framework (adapted from [34]).
Based on these ideas, we have developed HO3-SLAM (Human Object Occlusion Ordering SLAM) [34] (Fig. 10), where HOI boundary points are treated as key landmark feature points for mapping and localization, with a particular focus on loop closure detection. Since this occlusion boundary represents interactions between a dynamic human and a static occluder, it combines the desirable properties of both (Fig. 4): it is stably detectable via human tracking in both RGB and T domains and is invariant in appearance, geometry, and semantics—thus serving as a reliable landmark.
Figure 4: Human-Object Interaction (HOI) Features. Left: A rough sketch of the view, showing the human region (blue) and occluder objects (red). Right: The detected human region (white) and HOI feature points (red) in the thermal image. Although the occluder objects are barely visible in the thermal image, their characteristics can still be inferred from the arrangement of the boundary features.
III-C Key Difference from HO3-SLAM
The primary difference in D&D-SLAM compared to HO3-SLAM is the replacement of RGB cameras with thermal (T) cameras. This change introduces the following implications:
Human sensing with T cameras shows significant degradation in tracking and identification performance (Fig. 5) [35].
Static object sensing with T cameras is challenged by noise, ambient temperature effects, and low spatial resolution, as widely noted in thermal SLAM literature [7].
The idea of using humans as cues for mapping, as well as the concept of mapping dynamic objects, is not entirely novel—prior work [36] and [37] has explored these topics. However, this perspective has been scarcely explored within the context of HO3-SLAM or the broader human-as-landmark paradigm.
III-D Summary and Outlook
In summary, D&D-SLAM is a new SLAM paradigm characterized by dark scenes and dynamic landmarks, in contrast to conventional SLAM approaches designed for well-lit and static landmarks. We emphasize the urgent need for a new class of Image Change Pattern Recognition (ICPR) techniques that can operate under Non-Instrumented, Non-Color, Non-Active(-Electromagnetic) (NICA)  [^1] Instrumented: using external devices such as GPS sensors, wearable sensors attached to objects or people, or multi-view sensing systems like stereo cameras. Color: refers to sensors that measure surface reflectivity in the visible or near-visible spectrum, such as RGB cameras, grayscale cameras, and multispectral sensors. Active: refers to active(-electromagnetic) sensors such as LiDAR, ToF, structured light, etc. sensor modalities. The task of ICPR using NICA sensors in non-static and non-bright scenes has been explored in our previous work [31], but the exploration of D&D-SLAMis presented here for the first time. Thermal cameras offer a unique advantage—functioning in total darkness—but are inherently difficult to use, highlighting the “no free lunch (NFL)” principle. D&D-SLAM demands innovation across the ICPR pipeline: image change detection, counting, tracking, localization, and (re-)identification.
Figure 5: Human sensing capabilities of major passive sensors (and active sonar [31]), following the human sensing surveys in [38] and [35].
IV RGB-T Cooperative MPT
IV-A Formulation of MPT
@@ IV-C Cooperative Strategy
$\hat{y}_{t}=f_{\theta^{\text{RGB}}\theta^{\text{T}}}(x_{t}^{\text{T}})$ (5)
where $f_{\theta^{\text{RGB}}\theta^{\text{T}}}$ represents the fusion function integrating both models. This fusion model integrates the tracking results of the RGB teacher model and the T student model, selecting the result with the higher confidence score in each frame to achieve robust tracking. For each $i$ $(i\in[1,I^{t}])$, $y_{t}[i]=f^{\text{T}}(x_{t}^{\text{T}})$, $\tilde{y}_{t}[j]=f^{\text{RGB}}(x_{t}^{\text{T}})$, and for the corresponding pair $j$ $(j\in[1,I^{t}])$ satisfying $y_{t}^{\text{ID}}[i]=\tilde{y}_{t}^{\text{ID}}[j]$, the final fusion result is determined as follows:
@@ V Evaluation Experiment
V-A Multi-Person Tracking
Each MPT tracking method is evaluated using Multiple Object Tracking Accuracy (MOTA) and ID F1 score (IDF1) to assess tracking performance, along with False Positive (FP), True Positive (TP), and False Negative (FN) to evaluate false detections [46]. ID Switches (IDSW) and Higher Order Tracking Accuracy (HOTA) measure trajectory reconstruction, with HOTA providing a global tracking assessment. Unlike MOTA, HOTA balances precision and recall by decomposing detection, assignment, and localization for a more detailed analysis (Table II). These metrics gauge how effectively the proposed method tackles key challenges in Multiple Person Tracking (MPT): ID switching, detection failures, occlusion, and handling the appearance/disappearance of individuals. ID switching misassigns identities, detection failures disrupt continuity, and occlusion hides targets. Additionally, newly appearing or disappearing persons may be misidentified, reducing accuracy.
TABLE II: Tracking Accuracy
@@ V-A Multi-Person Tracking
BRIGHT 99.8 100.0 5 8273 3
RGB 26.3 42.6 33 2007 5371
T 89.4 95.0 631 8053 225
FUSION 87.9 94.8 661 8058 220
BRIGHT: $f_{\theta^{\text{RGB}}}(x_{\text{Bright}}^{\text{T}})$, RGB: $f_{\theta^{\text{RGB}}}(x_{\text{Dark}}^{\text{T}})$, T: $f_{\theta^{\text{T}}}(x_{\text{Dark}}^{\text{T}})$, FUSION: $f_{\theta^{\text{RGB}}\theta^{\text{T}}}(x_{\text{Dark}}^{\text{T}},x_{\text{Dark}}^{\text{T}})$.
TABLE III: Trajectory Prediction Performance
@@
V-B Loop Closure Detection
In a crowded and dynamic research lab, we conducted a loop closure experiment using a wheeled mobile robot equipped with RGB-T cooperative vision, similar to the setup shown in Figure 1. The robot utilized HO3 landmarks during the experiment. During this experiment, the robot moved and turned, capturing approximately 500 frames, corresponding to 2-3 minutes. Additionally, multiple humans were present in the scene, with some sitting at desks and working on desktop PCs, while others were seated on a sofa relaxing. In other words, they did not move much. Nevertheless, since the robot was in motion, a challenging scenario arose where multiple humans intersected within the scene.
The research question here is: “Can HO3 landmarks, useful for loop closing, be extracted from thermal vision?” We designed a simple image processing algorithm, and despite its simplicity, we found that it is still effective for loop closing.
The loop closure detection method is described below. In the explanation, the x-axis in the image points to the right, and the y-axis points downward. For each image frame:
  • Detect the human regions in the thermal image.
  • Detect and track human bounding boxes, assigning person IDs.
  • Obtain human regions with person IDs by performing pixel-wise AND operations between the bounding box and the human region.
For each person ID:
  • For each x-coordinate within the bounding box, find the pixel belonging to the human region that has the maximum y-coordinate.
This set of lower endpoints is referred to as landmarks, which are then recorded in the map along with their frame IDs. The current frame’s landmarks are matched to past landmarks in the map to detect loop closures:
  • Specifically, given a pair of current and past frames, the matching score from the current frame’s landmarks to past landmarks is searched and scored using RANSAC under an affine transformation, then ranked.
The performance of loop closure detection was evaluated based on top-5 accuracy. Specifically, in the RANSAC scoring ranking, if the ground-truth match ranked within the top 5, it was considered a success. The frequency of success was then measured and used as the performance metric. The ground-truth frame ID is defined as the frame ID of the viewpoint closest to the robot’s trajectory on the viewpoint coordinates.
Performance results are shown in Figure 8. A meaningful and high loop closure detection performance is achieved. In Figure 9 (b) left, two humans are severely occluded by boxes or partitions. Even so, the HO3 landmarks at the boundary between humans and occluders remain highly invariant, contributing to successful loop closure detection. As shown in Figure 9 (b) middle, humans that are not occluded have most of their lower endpoints lying within the bounding box, and in such cases, inappropriate landmarks can be excluded. Figure 9 (b) right shows that the other two HO3 landmarks function as useful landmarks. In this experiment, we demonstrated that HO3 landmarks are effective even compared to the best-known classical change detection and feature extraction methods. Our immediate challenge is to enhance and improve the robustness and accuracy of HO3 landmarks through the introduction of machine learning.
Figure 8: Loop closure detection performance. Vertical axis: Top-5 accuracy $(X/5,X\in\{0,1,2,3,4,5\})$. Horizontal axis: Sorted query image ID.
Figure 9: Loop closing experiments. (a) A query image (top), and a rough sketch of a looped trajectory with three viewpoints along the trajectory (bottom). (b) Thermal images (top) and rough scene sketches (bottom) acquired at those three viewpoints.
Figure 10: Block diagram of the HO3-SLAM framework (adapted from [34]).
V-C Thermal-Depth Cooperative Tracking
This section evaluates the person tracking performance based on cooperative vision using Thermal (TIR) and Depth images without relying on RGB images. The significance of this experiment lies in extending the application scope beyond RGB-T cooperative vision to encompass knowledge transfer among robots equipped with diverse sensor modalities. Such efforts represent an important first step toward realizing cross-robot SLAM. Comparisons with the RGB-T cooperative vision in the previous section are also conducted to examine the interchangeability of sensor modalities.
The TIR images were acquired by a HIKMICRO Pocket2 thermal camera. Depth images were generated by extracting a narrow field-of-view region corresponding to that of the TIR camera from 3D point clouds captured by a Velodyne VLP-16 LiDAR sensor, and projecting them into image form.
The dataset used in this experiment consists of:
  • Training set: 6,482 synchronized TIR-Depth image pairs (image resolution: 1440×1080)
  • Test set: A sequence of 3,797 depth image frames (image resolution: 640×430)
All images were manually annotated with person IDs and bounding boxes.
Figure 11 shows paired TIR and depth images. The depth images used here are not derived from high-precision ToF sensors but rather from a very low-resolution subset of a wide-area monitoring 3D LiDAR sensor (Velodyne VLP-16). As a result, fine details such as facial contours are mostly absent, making person identification extremely difficult for human observers.
In fact, many subjects found it easy to distinguish individuals in the TIR images but challenging in the depth images. As illustrated, facial shapes and contours are lost in the depth images.
Nevertheless, our study demonstrates that through knowledge transfer from TIR vision to depth vision, the appearance representation of each person can be effectively complemented, enabling highly accurate person tracking even with low-resolution depth images. This indicates that cooperative vision combined with learning-based knowledge transfer can significantly improve tracking performance in visually challenging environments.
Table IV presents the performance of the proposed depth-image-based D tracker. In addition to MOTA and IDF1, metrics such as HOTA and ID switch count also show favorable results, confirming the reliable person tracking performance achieved.
Figure 11: Cooperative tracking with TIR and depth vision. Top row: TIR images, Bottom row: Depth images.
TABLE IV: Person tracking performance with Thermal-Depth cooperative vision (D tracker)
MOTA IDF1 TP FP FN HOTA IDSW
95.6 98.0 10715 90 350 96.1 48
VI Conclusion and Future Works
In this study, we addressed the cooperative multi-person tracking (MPT) problem using RGB and thermal (T) cameras placed in the same viewpoint — for the first time. As expected, the RGB tracker completely failed under the Dark scene’s RGB modality and also exhibited degraded performance in thermal imagery. To address this issue, we confirmed that training the T tracker with pseudo-annotations generated by the RGB tracker is highly effective, achieving state-of-the-art performance in both Bright and Dark scenarios. We also explored a fusion-based tracker that integrates both RGB and T modalities; surprisingly, however, its performance ranked second. This suggests that a cooperative strategy — dynamically switching between RGB and T trackers depending on the situation — is more effective than a straightforward fusion. As a demonstration of applicability, we extend our previous work, Human-Object Occlusion Ordering (HO3)-SLAM, to thermal (T) cameras by leveraging the proposed thermal MPT tracker. In the experimental section, we present a unique proof-of-concept loop-closure detection experiment in which the human-object-interaction (HOI) between humans and their occluders are treated as pseudo—dynamic landmark feature points (albeit with frequent false negatives). Whereas conventional SLAM systems typically focus on mapping static landmarks under well-lit conditions, our approach takes a first step toward a new paradigm — Dynamic-Dark SLAM — which aims to enable robust mapping across the full spectrum of landmark dynamics (static and dynamic) and lighting conditions (illumination and complete darkness).